About what percent of products last 15 more days


1
Expert's answer
2021-10-25T14:44:53-0400

Let us use the following example:

The shelf life of a particular dairy product is normally distributed with a mean of 12 days and a standard deviation of 3 days. About what percent of the products last 15 or more days?

"\mu = 12 \\

\sigma = 3 \\

P(X≥15) = 1 -P(X<15) \\

= 1 -P(Z < \frac{15-12}{3}) \\

= 1 -P(Z< 1) \\

= 1 -0.8413 \\

= 0.1587"